OpenStudy (anonymous):

-7x - y = -6 x - 2y = -12 pleaaase help? :)

OpenStudy (anonymous):

x=0, y=6

OpenStudy (anonymous):

can i know how did you do it? :)

OpenStudy (anonymous):

isolate y for -7x - y = -6 y=-7x+6 sub y=-7x+6 into x - 2y = -12 x - 2(-7x-6) = -12 solve for x, you should get 0 plug x=0 back in y=-7x+6 solve for y, you should get y=6

OpenStudy (anonymous):

thank you!!!!!!!! :)